# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

### v2

@@ 5 Numerical experiments
Figure 3: Performance of RU-MLMC, GRR-MLMC and TGRR-MLMC A. Two-moon, B. Lotka-Volterra C. M/G/1 queue model, blue, green, and red correspond to RU-MLMC, GRR-MLMC, and TGRR-MLMC respectively.
We present our results for the baseline models in Figure 2 and Figure 3. It is observed that both unbiased methods are inferior to the biased ones in some cases. When comparing RU-MLMC with GRR-MLMC in the case of the Lotka-Volterra and M/G/1 queue models, we find that the unbiased method can greatly benefit from variance reduction. In the case of Lotka-Volterra, when comparing TGRR and GRR methods, we conclude that the gradient information $\Delta\rho_{\phi,\ell}$ where $\ell\geq\overline{m}$ not only does not contribute to the overall performance but also make it worse when measured with C2ST. As suggested by Theorem 4.2, this could be due to the domination of variance, where the effect of reducing variance is superior to reducing bias.
In conclusion, in the case of SNPE where the complex model of $q_{F(x,\phi)}(\theta)$ is used for density estimation, variance tends to dominate the bias as it is suggested by Theorem 4.2, indicating that excessive variance could seriously affect the training process of the density estimator. Instead of seeking unbiasedness, one should try to strike a balance between bias, average computational cost, and variance of the gradient in this case.
It is worth noting that in the case of the Two-moon problem, RU-MLMC method delivers comparable performance to the other two methods with lower computational cost and time requirements. Rather than employing any of these methods indiscriminately, it is crucial to undertake a comprehensive and thoughtful analysis of the specific problem at hand.
In Table 2, we also compare TGRR-MLMC with sequential neural score estimation (SNSE) proposed by [54], which is a state-of-the-art likelihood-free inference. We find that TGRR-MLMC demonstrates consistent improvement over SNSE across three baseline tasks. As shown in Table 3, TGRR-MLMC takes approximately 1.2–1.5 fold computational time of SNSE. However, SNSE consumes up to 60GB of GPU memory for a batch size of 128, while our method requires only 5GB for a batch size of 500, which demonstrates a substantial advantage in memory efficiency. For the hyperparameter of SNSE, we mainly follow the default setting from [13] and change the maximum iteration to 1,000 and samples per round to 1,000 to align with our settings.
To investigate the performance of the proposed MLMC method scaling with higher dimensional problems, we infer the single-compartment Hodgkin-Huxley type neuron model [49, 23] with 8 parameters of interest. The results are deferred to Appendix E.
@@ 6 Concluding remarks
The atomic APT method is a heuristic nested estimation algorithm. Given that its algorithm design is not friendly to theoretical analysis, this paper first proposes a nested APT method with similar performance, thus enabling a series of convergence analyses to quantify the impact of inner and outer sample sizes on the computational complexity. Given the biased nature of the nested APT method, this paper conducts a comprehensive exploration based on a series of unbiased MLMC to eliminate the bias in the nested APT method. This paper also proposes several theorems to quantify the bias, variance, and average computational cost of the loss function and its gradient estimators. Additionally, since the optimization method based on SGD is used to update the neural network parameters, it is well-known that unbiased methods suffer from excessively high variance in gradient estimation. To address this issue, this paper proposes using several variants of the truncated MLMC method to reduce the variance. A series of numerical experiments verify the effectiveness of these variants.
We acknowledged some limitations regarding the proposed MLMC methods. The methods may suffer from excessive variance accompanied by high dimensional problems and complex neural network structures, which lead to inferior performances in some extent. A possible way to reduce the variances of MLMC estimators is to use quasi-Monte Carlo (QMC) methods, which use low discrepancy (more balanced) points to generate samples from the proposal. QMC-based quadrature can yield a faster convergence rate than plain Monte Carlo, and its application in nested MLMC methods have already been investigated in [29]. Secondly, MLMC takes more computational time than atomic APT due to the gradient calculation at multiple levels. How to speed up the algorithm of MLMC is left for future work.
@@
[…]
@@
$+\left\|\frac{\nabla Z_{\phi}(x,\phi)}{Z(x,\phi)}\right\|_{2}\left((S^{(a)}_{\ell-1})^{2}+(S^{(b)}_{\ell-1})^{2}+2(S_{\ell})^{2}\right).$ (50)
@@ Appendix B Normalizing flows as a density estimator
Therefore, it is determined by the bijection $f$ and its inverse $f^{-1}$, which are often implemented by composing a series of invertible neural-network modules. Neural spline flows [13] used in our paper is built upon coupling layers, it maps $\mathbf{x}$ to $\mathbf{y}$ by using monotonic rational-quadratic splines, where each interval is defined by a monotonic rational-quadratic function taking the form of a ratio of two quadratic polynomials. These functions are easy to differentiate and, when restricted to monotonic segments, they are also analytically invertible. Moreover, both the derivative and the height at each knot are strictly more flexible. The spline itself maps the interval $[-B,B]$ to $[-B,B]$, and outside this range, the transformation is defined as the identity, resulting in a linear ‘tail’ that allows the overall transformation to accept unconstrained inputs.
@@
We specify the spline using $K-1$ arbitrary positive derivative values $\{\delta^{(k)}\}_{k=1}^{K-1}$ at the $K-1$ interior knots $(u^{(k)},v^{(k)})_{k=1}^{K-1}$, and set the boundary derivatives to $\delta^{(0)}=\delta^{(K)}=1$ in order to meet the linear tails.
@@
    A neural network is used to take $\mathbf{x}_{1:D^{\prime}-1}$ as input and output an unconstrained parameter vector $\phi_{i}$ of length $13K-$ 1 for each $i=D^{\prime},\ldots,D$.
@@
    Compute $\mathbf{y}_{i}=\Gamma_{\phi_{i}}(\mathbf{x}_{i})$ in parallel for $i=D^{\prime},\dots,D$, where $\Gamma_{\phi_{i}}$ is an invertible function parameterized by $\phi_{i}=[\phi_{i}^{w},\phi_{i}^{h},\phi_{i}^{d}]$, where $\phi_{i}^{w}$ and $\phi_{i}^{h}$ are of length $K$, and $\phi_{i}^{d}$ is of length $K-1$, vectors $\phi_{i}^{w}$ and $\phi_{i}^{h}$ are used to obtain the widths and heights of $K$ intervals. The cumulative sums of these widths and heights define the $K+1$ knot points $\{(u^{(k)},v^{(k)})\}_{k=0}^{K}$. Vector $\phi_{i}^{d}$ is obtained the derivative values $\{\delta^{(k)}\}_{k=1}^{K-1}$ at the interior knots.
@@
Figure 4: Approximated posterior for nested APT and atomic APT. A. Two-moon model, from left to right: available ground truth, atomic APT with inner samples $M=100$, nested APT with inner samples $M=100$. B. Lotka-Volterra model, from left to right: ground truth simulated with SMC-ABC [4], atomic APT with $M=100$, nested APT with $M=100$. C. M/G/1 queue model, the setting are the same with Lotka-Volterra. We have plotted the scatter plot for each of the 2D sub-spaces. The histogram for each $\theta_{i},i=1,\dots,4$ is plotted on the diagonal with the ground truth parameter marked with dotted lines.

### v3

@@ 5 Numerical experiments
Figure 3: Performance of RU-MLMC, GRR-MLMC and TGRR-MLMC A. Two-Moon, B. Lotka-Volterra C. M/G/1 queue model, blue, green, and red correspond to RU-MLMC, GRR-MLMC, and TGRR-MLMC, respectively.
We present our results for the baseline models in Figure 2 and Figure 3. We observe that both unbiased methods are inferior to the biased ones in some cases. When comparing RU-MLMC with GRR-MLMC in the case of the Lotka-Volterra and M/G/1 queue models, we find that the unbiased method can greatly benefit from variance reduction. In the case of Lotka-Volterra, when comparing TGRR and GRR methods, we conclude that the gradient information $\Delta\rho_{\phi,\ell}$ where $\ell\geq\overline{m}$ not only does not contribute to the overall performance but also makes it worse when measured with C2ST. As suggested by Theorem 4.2, this could be due to the domination of variance, where the effect of reducing variance is superior to reducing bias.
In the case of SNPE, where the complex model of $q_{F(x,\phi)}(\theta)$ is used for density estimation, variance tends to dominate the bias as it is suggested by Theorem 4.2, indicating that excessive variance could seriously affect the training process of the density estimator. Instead of seeking unbiasedness, one should try to strike a balance between bias, average computational cost, and variance of the gradient in this case.
Taking these empirical trends and the bias-variance trade-off, we conclude the following practical guidance on choosing an appropriate MLMC variant.
  1. 1.
    Under tight computational budgets or complex tasks, TGRR-MLMC often yields the most stable training.
  2. 2.
    When unbiasedness is required and the effect of variance is moderate, GRR-MLMC is a favorable option.
  3. 3.
    When both unbiasedness and minimal cost are priorities, RU-MLMC is preferable.
In our experiments, this pattern is reflected across tasks. It suffices to use RU-MLMC for the simple Two-Moon model. GRR-MLMC performs well on the intermediate M/G/1 queue model. TGRR-MLMC is a better choice for the more challenging Lotka–Volterra model.
In Table 2, we also compare TGRR-MLMC with sequential neural score estimation (SNSE) proposed by [54], which is a state-of-the-art likelihood-free inference method. We find that TGRR-MLMC achieves a lower mean C2ST than SNSE on all three tasks; however, it requires 1.2–1.5× wall-clock time, as shown in Table 3. Notably, SNSE consumes up to 60 GB of GPU memory for a batch size of 128, whereas our method requires only 5 GB for a batch size of 500, demonstrating a substantial advantage in memory efficiency. For the hyperparameter of SNSE, we mainly follow the default setting from [13] and change the maximum iteration to 1,000 and samples per round to 1,000 to align with our settings.
To investigate how the proposed MLMC method scales to higher-dimensional problems, we infer the single-compartment Hodgkin-Huxley type neuron model [49, 23] with 8 parameters of interest. The results are deferred to Appendix E.
@@ 6 Concluding remarks
We revisited APT by replacing the atomic heuristic with a nested formulation to allow rigorous convergence analysis of learning. We explore a family of MLMC nested estimators to accommodate different scenarios, as lower computational complexity alternatives to the single-level nested estimator. Our theory quantifies the orders of bias, variance, and average cost for the loss and gradient estimators. Numerical experiments confirm the effectiveness of the proposed methods for various tasks.
We acknowledge some limitations regarding the proposed methods. Firstly, they may suffer from excessive variance accompanied by high-dimensional problems and complex neural network structures, which lead to inferior performance to some extent. A possible way to reduce the variances of MLMC estimators is to use quasi-Monte Carlo (QMC) methods, which use low-discrepancy (intuitively, more uniform) points to generate samples from the proposal. QMC-based quadrature can yield a faster convergence rate than single-level Monte Carlo, and its application in nested MLMC methods has already been investigated in [29]. Secondly, MLMC takes more computational time than atomic APT due to the gradient calculation at multiple levels. How to speed up the algorithm of MLMC is left for future work.
@@
Let $\{a_{i}\}_{i=1}^{d}$ be any real finite sequence and $\{b_{i}\}_{i=1}^{d}$ be a positive finite sequence, for any $j>0$ we have $a_{j}\leq b_{j}\ \underset{i}{\max}\left\{{a_{i}}/{b_{i}}\right\}$. Taking summation on both sides yields
@@
$+\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2}\left((S^{(a)}_{\ell-1})^{2}+(S^{(b)}_{\ell-1})^{2}+2(S_{\ell})^{2}\right).$ (50)
@@ Appendix B Normalizing flows as a density estimator
Therefore, it is determined by the bijection $f$ and its inverse $f^{-1}$, which are often implemented by composing a series of invertible neural-network modules. Neural spline flows [13] used in our paper is built upon coupling layers, it maps $\mathbf{x}$ to $\mathbf{y}$ by using monotonic rational-quadratic splines, where each interval is defined by a monotonic rational-quadratic function taking the form of a ratio of two quadratic polynomials. These functions are easy to differentiate and, when restricted to monotonic segments, they are also analytically invertible. Moreover, both the derivative and the height at each knot are strictly more flexible. The spline itself maps the interval [B,B][-B,B] to [B,B][-B,B], and outside this range, the transformation is defined as the identity, resulting in a linear ‘tail’ that allows the overall transformation to accept unconstrained inputs.
@@
We specify the spline using K1K-1 arbitrary positive derivative values $\{\delta^{(k)}\}_{k=1}^{K-1}$ at the K1K-1 interior knots $(u^{(k)},v^{(k)})_{k=1}^{K-1}$, and set the boundary derivatives to $\delta^{(0)}=\delta^{(K)}=1$ in order to meet the linear tails.
@@
    A neural network is used to take $\mathbf{x}_{1:D^{\prime}-1}$ as input and output an unconstrained parameter vector $\phi_{i}$ of length 3K13K-1 for each $i=D^{\prime},\ldots,D$.
@@
    Compute $\mathbf{y}_{i}=\Gamma_{\phi_{i}}(\mathbf{x}_{i})$ in parallel for $i=D^{\prime},\dots,D$, where $\Gamma_{\phi_{i}}$ is an invertible function parameterized by $\phi_{i}=[\phi_{i}^{w},\phi_{i}^{h},\phi_{i}^{d}]$, where $\phi_{i}^{w}$ and $\phi_{i}^{h}$ are of length $K$, and $\phi_{i}^{d}$ is of length K1K-1, vectors $\phi_{i}^{w}$ and $\phi_{i}^{h}$ are used to obtain the widths and heights of $K$ intervals. The cumulative sums of these widths and heights define the $K+1$ knot points $\{(u^{(k)},v^{(k)})\}_{k=0}^{K}$. Vector $\phi_{i}^{d}$ is obtained the derivative values $\{\delta^{(k)}\}_{k=1}^{K-1}$ at the interior knots.
@@
Figure 4: Approximated posterior for nested APT and atomic APT. A. Two-Moon model, from left to right: available ground truth, atomic APT with inner samples $M=100$, nested APT with inner samples $M=100$. B. Lotka-Volterra model, from left to right: ground truth simulated with SMC-ABC [4], atomic APT with $M=100$, nested APT with $M=100$. C. M/G/1 queue model, the settings are the same as Lotka-Volterra. We have plotted the scatter plot for each of the 2D subspaces. The histogram for each $\theta_{i},i=1,\dots,4$ is plotted on the diagonal with the ground truth parameter marked with dotted lines.